OpenStudy (anonymous):

Question with log, solve for x: (log2X+logX)/(log2X-logX) = 3 (base of log is the same)

OpenStudy (anonymous):

\[\frac{ \log2x+logx }{ \log2x-logx } = 3\]

OpenStudy (anonymous):

show me the steps please

OpenStudy (anonymous):

X=2

OpenStudy (anonymous):

Sorry .

OpenStudy (anonymous):

yes x=2

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

log2x +logx = log2x^2

OpenStudy (kropot72):

From the rules of logs \[\log_{} 2x+\log_{} x=\log_{} 2x ^{2}\] Do you understand this step?

OpenStudy (anonymous):

that's how i translated it for the numerator yes

OpenStudy (anonymous):

for the denominator it's log2x/x

OpenStudy (anonymous):

so it's log2

OpenStudy (kropot72):

So othe denominator is log 2. Correct.

OpenStudy (anonymous):

and then???

OpenStudy (kropot72):

So we now have by cross multiplying \[\log_{} 2x ^{2}=3\log_{} 2\] Understood?

OpenStudy (anonymous):

oh i see!

OpenStudy (anonymous):

so what i did was

OpenStudy (kropot72):

and \[3\log_{} 2=\log_{} 2^{3}=\log_{} 8\]

OpenStudy (anonymous):

log2x^2/log2 , i just cancelled out the 2's on top and bottom, leaving logx^2=3, so that's wrong

OpenStudy (anonymous):

yes i understand what you did there it makes sense, i just want to understand why i'm wrong

OpenStudy (kropot72):

Yes, that was wrong. \[\log_{} 2x ^{2}=\log_{} 8\] So\[2x ^{2}=8\]

OpenStudy (anonymous):

i just have one quick question, when i see this \[\frac{ \log2x ^{2} }{ \log2? }\] is it equivalent to log2x^2-log2 ?

OpenStudy (anonymous):

(take out the question mark beside log2)

OpenStudy (kropot72):

To understand where you went wrong, please look at this example: \[\frac{\log_{} (2 \times 100)}{\log_{} 2}=\frac{\log_{} 200}{\log_{} 2}=7.644\] If you try your 'cancelling' you would get a result of 2, which is obviously wrong. Does this help you?

OpenStudy (anonymous):

hmm interesting. it makes more sense now

OpenStudy (anonymous):

yeah it's really more clear when you see it with actual numbers thank you so much.

OpenStudy (kropot72):

The answer to your quick question is no, the reason being that both terms are already logarithms. Therefore you do not apply the rule that you would apply to a quotient of real numbers.

OpenStudy (anonymous):

thank you sir i'll try to chew on it more, it was very helpful.

OpenStudy (kropot72):

I'm glad to be of help :)